1. What is the present value of $1 to be paid out in seven years with an interest rate of 10%? (2 pts)
  2. What is the present value of $80 to be paid out in seven years with an interest rate of 10% (2 pts)
  3. What is the present value of $400,000 to be paid out in ten years with an interest rate of 6%? (2 pts)
  4. What is the future value of $100,000 in five years at an annual interest rate of 5%? (2 pts)
  5. The following cash flow is from an ROI and represents the net income expected over 5 years. Assuming a 10% cost of capital, what is the Net Present Value? (5 pts)

                 (in thousands 4

   6. What is the payback period for the question above? (2 pts)
Answer #1

(A) Present Value of $ 1 to be paid out after 7 years at 10% = 1 / (1.1)^(7) = $ 0.513

(B) Present Value of $ 80 to be paid out after 7 years at 10% = 80 / (1.1)^(7) = $ 41.053

(C) Present Value of $ 400000 to be paid out after 10 years at 6% = 400000 / (1.06)^(10) = $ 223357.91

(D) Future Value of $ 100000 in five years at an annual rate of 5 % = 100000 x (1.05)^(5) = $ 127628.16

NOTE: Please raise separate queries for solutions to the remaining unrelated questions as one query is restricted to the solution of only one question with a maximum of four sub-parts.